# SONICS: Synthetic Or Not - Identifying Counterfeit Songs

### v2

@@ Abstract
The recent surge in AI-generated songs presents exciting possibilities and challenges. While these tools democratize music creation, they also necessitate the ability to distinguish between human-composed and AI-generated songs for safeguarding artistic integrity and content curation. Existing research and datasets in fake song detection only focus on singing voice deepfake detection (SVDD), where the vocals are AI-generated but the instrumental music is sourced from real songs. However, this approach is inadequate for contemporary end-to-end AI-generated songs where all components (vocals, lyrics, music, and style) could be AI-generated. Additionally, existing datasets lack lyrics-music diversity, long-duration songs, and open fake songs. To address these gaps, we introduce SONICS, a novel dataset for end-to-end Synthetic Song Detection (SSD), comprising over 97k songs with over 49k synthetic songs from popular platforms like Suno and Udio. Furthermore, we highlight the importance of modeling long-range temporal dependencies in songs for effective authenticity detection, an aspect overlooked in existing methods. To capture these patterns, we propose a novel model, SpecTTTra, that is up to 3 times faster and 6 times more memory efficient compared to popular CNN and Transformer-based models while maintaining competitive performance. Finally, we offer both AI-based and Human evaluation benchmarks, addressing another deficiency in current research.
**footnotetext: Equal contribution22footnotetext: Code and Data will be available at https://github.com/awsaf49/sonics
@@ 1 Introduction
The rapid advancements in AI-generated music present a substantial threat to the music industry, potentially reducing the demand for professional musicians and stifling new talent development [19, 5]. To preserve the unique value of human creativity, it is crucial to develop robust methods for detecting AI-generated music, ensuring a fair and vibrant creative ecosystem.
@@
Singing Voice Synthesis (SVS) [16] and Singing Voice Conversion (SVC) [11] have recently achieved significant progress, enabling the creation of synthetic singing voices that closely mimic real singers’ styles. When combined with instrumental music from real songs, these synthetic voices can produce convincing counterfeit songs. Although related to synthetic speech detection, detecting fake songs is particularly challenging due to the unique rhythmic patterns and artistic vocal traits of singing. To address this, researchers have turned their attention to Singing Voice Deepfake Detection (SVDD) [30, 34, 33]. However, current methods relying on datasets composed of SVS and SVC-generated songs face several limitations. These datasets are bound to use instrumental music from real songs, leading to artifacts like the "Karaoke effect" (volume discrepancies between music and vocals) and limited music-lyrics diversity. Moreover, existing methods often overlook the long-context temporal relationships inherent in songs, such as repeated verses, music, rhythm, and emotional dynamics, which are critical for effective detection. The short duration of songs in current datasets further hampers the use of these patterns. Additionally, copyright restrictions on some existing datasets limit the public availability of generated fake songs, hindering broader usage. Furthermore, the SVDD task requires separate tools for voice identification and separation during data processing [30, 34], increasing computational overhead.
@@
Recently, platforms like Suno [24] and Udio [6] have gained significant traction on social media. They can synthesize not only vocals but also entire songs, including synthetic music, styles, and lyrics, further complicating the situation. Due to their end-to-end nature, these fake songs differ significantly from those generated by SVS and SVC methods, rendering existing SVDD methods and datasets inadequate for detecting them. This necessitates an urgent need for a detection system capable of identifying these end-to-end fake songs.
@@
To address these shortcomings, we introduce SONICS, a large-scale dataset comprising over 97k songs, including more than 49k end-to-end synthetic songs generated by Suno and Udio, alongside 48k real songs collected from YouTube. With an average duration of 145 seconds, the SONICS dataset supports the use of long-context relationships in songs for accurate fake song detection. Our dataset also addresses the music-lyrics diversity issue by including fake songs generated with both real and synthetic lyrics and music. Additionally, the fake songs in our dataset are free from copyright issues, as they are generated using paid subscriptions. A unique feature of SONICS is that it includes the text lyrics of songs, which can aid future research.
@@
Despite the availability of long songs in our dataset, utilizing long-context relationships presents additional challenges. For instance, CNN-based models struggle to capture long-range dependencies due to their inherently local receptive fields. While Transformer-based models can capture these dependencies with global attention, they are computationally expensive with longer audio inputs. To mitigate this trade-off, we introduce Spectro-Temporal Tokens Transformer (SpecTTTra), which uses a Spectro-Temporal Tokenizer to significantly reduce computational costs while enjoying global attention. Our contributions are summarized as follows:
@@
  • We introduce SONICS, a large-scale dataset for synthetic song detection that addresses the limitations of existing datasets, including the absence of end-to-end fake songs, limited music-lyrics diversity, short-duration songs, and open fake songs.
@@
  • We propose the SpecTTTra model, which effectively captures long-context temporal relationships in songs at a low computational cost.
  • We offer a human benchmark for the fake song detection task, a feature missing in previous works.
  • We also establish a traditional AI benchmark, evaluating models across data partitions for deeper insights.
@@
CtrSVDD Multi (no En.) 4.87 14 164 32,312 188,486 220,798
@@
Table 1: Comparison of Proposed and Existing Fake Song Datasets
@@ 2 Related Works
Synthetic Speech Detection: The domain of synthetic speech detection, closely tied to synthetic song detection through their shared audio modality, has been extensively explored due to advancements in voice conversion [37] and synthesis techniques [28]. These advancements have spurred the development of audio spoofing attacks on speaker verification systems and deepfake audio targeting human listeners [13]. Countermeasures include Light CNN (LCNN) with Max-Feature-Map activations [15], Transformer encoders with ResNet architectures [36], RawNet2 with sinc layers and GRU blocks [25], and heterogeneous graph attention networks [12]. However, the unique complexities of songs—such as rhythm, melody, and emotional nuance—present challenges that traditional speech detection methods are not equipped to handle [30, 34]. Thus, following CtrSVDD [33], we opted not to conduct similar experiments in our study.
@@
Synthetic Song Detection: Synthetic song detection, a relatively newer and more complex challenge, has gained attention recently. In 2023, SingFake [34] introduced a dataset of counterfeit songs using Singing Voice Conversion (SVC), along with the task of Singing Voice Deepfake Detection (SVDD) and associated model benchmarks. Subsequent work [30, 33] combined Singing Voice Synthesis (SVS) with SVC to create phoneme-based songs, leading to specialized detection datasets. Methods in this area include convolutional networks for feature extraction followed by graph neural network classification [33], wav2vec2-based extraction coupled with graph neural networks [26], and Linear-Frequency Cepstral Coefficients (LFCC) used with ResNet18 models [35]. However, SVC and SVS-based datasets always retain original background music, leading to a detectable "Karaoke effect" artifact. Recent end-to-end fake songs by Suno [24] and Udio [6], which can produce divergent fake songs where all musical components (e.g., background music, styles, and lyrics) can be synthetic, presenting a severe detection challenge.
@@
Long Audio Classification: Songs exhibit long-range temporal patterns, such as repeated verses and rhythms, setting them apart from speech [1]. Despite their potential to enhance detection performance, these patterns have been largely overlooked in existing methods [30, 34, 33]. Meanwhile, long audio classification remains an uncharted area in audio research. Although automatic speech recognition handles long audio data [14], it struggles with end-to-end processing of extended audio due to its high computational cost, often using sliding window techniques [9, 21] to manage costs. This further complicates the task of leveraging long-context features for synthetic song detection.
@@ 3 Methodology
Split Label Source Count
Train Full Fake Suno-v3.5 774
Udio-130 656
Half Fake Suno-v3.5 4558
Mostly Fake Suno-v3.5 11863
Udio-130 16172
Real YouTube 32686
Total 66709
Test Full Fake Suno-v2 100
Suno-v3 206
Suno-v3.5 77
Udio-130 53
Udio-32 203
Half Fake Suno-v2 309
Suno-v3 634
Suno-v3.5 396
Mostly Fake Suno-v2 1350
Suno-v3 2797
Suno-v3.5 1117
Udio-130 1566
Udio-32 3970
Real YouTube 13237
Total 26015
Table 2: Dataset Summary
@@
3.1 SONICS Dataset
The development of a modern synthetic song detection system necessitates a dataset that meets several stringent criteria, which are conspicuously absent in existing music datasets. These criteria include: 1) songs where all components—not just vocals—can be AI-generated; 2) song lengths sufficient to capture long-term contextual relationships; 3) a diverse spectrum of music-lyrics combinations; and 4) a quantity of data substantial enough to serve as a generative model benchmark. Addressing these needs, we introduce the SONICS dataset, a comprehensive collection of end-to-end AI-generated songs produced using the latest audio generative models, spanning lengths from 32 to 240 seconds. The dataset encompasses an extensive array of music-lyrics styles, amassing a total of 97,164 songs. A detailed comparison of SONICS with existing datasets is presented in Table 1, clearly illustrating that datasets such as FSD [30], SingFake [34], and CtrSVDD [33] fall short of fulfilling all the outlined criteria when juxtaposed with SONICS. Additionally, a comprehensive distribution summary of the SONICS dataset is provided in Table 2.
3.1.1 Real Songs
The dataset’s Real songs segment comprises original compositions by human artists. This portion of the dataset was initially compiled by random sampling from the Genius Lyrics Dataset [10], which provides metadata including lyrics, titles, and artist information. Subsequently, leveraging this metadata, an API-based dynamic search was conducted on YouTube to retrieve the corresponding audio files [2]. This process yielded 48,090 original songs performed by 9,096 artists.
3.1.2 Fake Songs
The Fake songs subset consists of compositions generated by AI models guided by two primary text inputs: 1) lyrics and 2) song-style (e.g., instruments, genre, vocal type). Based on the nature of these inputs, the Fake songs are categorized into three distinct groups: i) Full Fake (FF)—featuring both AI-generated lyrics and song-style; ii) Mostly Fake (MF)—where lyrics are AI-generated based on real lyrics features and song-style is derived from real songs; and iii) Half Fake (HF)—where the lyrics are directly sourced from real songs, with song-style also extracted from real compositions.
To generate FF songs, we curated a rich set of metadata, including 57 broad topics (e.g., friendship, betrayal), 292 specific topics (e.g., star trek, pokemon), 49 music genres (e.g., rock, metal), and 72 moods (e.g., calm, angry). Random combinations of these elements were used to generate final lyrics and styles via a Large Language Model (LLM). The complete pipeline for FF song generation is illustrated in Fig. 1. In contrast, the creation of MF and HF songs involved extracting song- styles from real songs using the multimodal LLM Gemini 1.5 Flash, which uniquely supports audio input. For MF songs, lyrics were generated using the GPT-4o LLM, which extracts features from existing real song lyrics to produce new lyrics with a similar distribution, avoiding direct replication of existing content. The pipeline for generating MF songs is detailed in Fig. 2. The HF songs, differing only in the use of real lyrics directly sourced from existing songs, follow a similar process. It is noteworthy that, Udio models cannot generate songs for lyrics from real songs; thus, HF songs contain only Suno models. Using this curated metadata, a variety of audio generative models were employed to produce a total of 49,074 Fake songs, comprising 6,132 HF, 40,769 MF, and 2,173 FF songs. The prompt templates used in LLM to extract lyrics feature and create lyrics along with song- styles can be found in the appendix.
For generating these end-to-end songs, we utilized the only two currently available audio generative model families: Suno [24] and Udio [6]. Specifically, we incorporated three variants from Suno—Suno v2, v3, and v3.5. Suno v2, the earliest iteration, generates songs up to 80 seconds (sec.) in length. Suno v3 extends this capability to 120 sec, offering enhanced audio quality and broader stylistic support. The latest iteration, Suno v3.5, produces songs up to 240 sec, featuring improved song structure and vocal flow. From the Udio family, we used Udio-32 and Udio-130, capable of generating songs of 32 and 130 sec, respectively. It is noteworthy that, both the Suno and Udio models generate two songs per (lyrics, style) input.
3.2 Model
The architecture of an audio classification model is influenced by how the input audio is processed for feature extraction. In our benchmark, we primarily use Spectrograms more specifically Mel-spectrogram due to their versatile usage and effective performance across various audio processing tasks [8, 21, 22, 20, 23]. Given the 2D image-like nature of Spectrogram, image classifier models are used for classifying the audio. Given that image classification models can broadly be divided into two categories: CNN and Transformer-based architectures, in our work, we employ popular models from both categories, such as ConvNeXt [18], ViT [7], for benchmarking. However, these models encounter challenges when dealing with the inherent long-range temporal dependencies in songs. Specifically, ConvNeXt struggles to capture long-range dependencies due to its local receptive field, a characteristic inherited from CNNs. On the other hand, while ViT is capable of capturing long-range dependencies, it incurs significant computational costs as the number of patches/tokens rapidly increases with longer audio inputs. To address this issue, we introduce SpecTTTra, which utilizes global attention similar to ViT but employs a Spectro-Temporal Tokenizer to reduce the number of tokens considerably, thereby lowering computational costs and enhancing efficiency significantly.
3.2.1 Spectro-Temporal Tokens Transformer

### v4

@@ Abstract
The recent surge in AI-generated songs presents exciting possibilities and challenges. These innovations necessitate the ability to distinguish between human-composed and synthetic songs to safeguard artistic integrity and protect human musical artistry. Existing research and datasets in fake song detection only focus on singing voice deepfake detection (SVDD), where the vocals are AI-generated but the instrumental music is sourced from real songs. However, these approaches are inadequate for detecting contemporary end-to-end artificial songs where all components (vocals, music, lyrics, and style) could be AI-generated. Additionally, existing datasets lack music-lyrics diversity, long-duration songs, and open-access fake songs. To address these gaps, we introduce SONICS [^1] Code & Data available at https://github.com/awsaf49/sonics, a novel dataset for end-to-end Synthetic Song Detection (SSD), comprising over 97k songs (4,751 hours) with over 49k synthetic songs from popular platforms like Suno and Udio. Furthermore, we highlight the importance of modeling long-range temporal dependencies in songs for effective authenticity detection, an aspect entirely overlooked in existing methods. To utilize long-range patterns, we introduce SpecTTTra, a novel architecture that significantly improves time and memory efficiency over conventional CNN and Transformer-based models. For long songs, our top-performing variant outperforms ViT by 8% in F1 score, is 38% faster, and uses 26% less memory, while also surpassing ConvNeXt with a 1% F1 score gain, 20% speed boost, and 67% memory reduction.
@@ 1 Introduction
The rapid advancements in AI-generated music present a substantial threat to the music industry, potentially reducing the demand for professional musicians and stifling new talent development (McMahon, 2024; Derbyshire et al., 2023). To preserve the unique value of human creativity, it is crucial to develop robust methods for detecting AI-generated music, ensuring a fair and vibrant creative ecosystem.
@@
Singing Voice Synthesis (SVS) (Liu et al., 2022a) and Singing Voice Conversion (SVC) (Jayashankar et al., 2023) have recently achieved significant progress, enabling the creation of synthetic singing voices that closely mimic real singers’ styles. When combined with instrumental music from real songs, these synthetic voices can produce convincing counterfeit songs. Although related to synthetic speech detection, detecting fake songs is particularly challenging due to the unique rhythmic patterns and artistic vocal traits of singing (Zang et al., 2024b). To address this, researchers have turned their attention to Singing Voice Deepfake Detection (SVDD) (Xie et al., 2024; Zang et al., 2024b; a). However, current methods relying on datasets composed of SVS and SVC-generated songs face several limitations. These datasets are bound to use only instrumental music from real songs, leading to artifacts like the “Karaoke effect” (volume discrepancies between music and vocals) and limited music-lyrics diversity. Moreover, existing methods overlook the long-context temporal relationships inherent in songs, such as repeated verses, music, rhythm, and emotional dynamics, which are critical for effective detection. Availability of only short duration songs in current datasets further hampers the use of these patterns. Additionally, copyright restrictions on some existing datasets limit the public availability of generated fake songs, hindering broader usage. Furthermore, the SVDD task requires separate tools for voice identification and separation during data processing (Xie et al., 2024; Zang et al., 2024b), increasing computational overhead.
@@
Recently, platforms like Suno [^2] https://suno.com, 2022. Accessed: 2024-06-27 and Udio [^3] https://udio.com, 2023. Accessed: 2024-07-09 have gained significant traction on social media. They can synthesize not only vocals but also entire songs, including synthetic music, styles, and lyrics, further complicating the situation. Due to their end-to-end nature, these fake songs differ significantly from those generated by SVS and SVC methods, rendering existing SVDD methods and datasets inadequate for detecting them. This necessitates an urgent need for a detection system specifically designed for end-to-end synthetic song detection (SSD).
@@
To address these shortcomings, we introduce SONICS, a large-scale dataset comprising 97,164 songs (4,751 hours), including 49,074 end-to-end synthetic songs (1,971 hours) generated by Suno and Udio, alongside 48,090 real songs (2,780 hours) curated from YouTube. With an average duration of 176 second (sec.), the SONICS dataset supports the use of long-context relationships in songs for accurate fake song detection. Furthermore, SONICS addresses the issue of music-lyrics diversity by including a wide range of music styles and both real and synthetic lyrics. A unique feature of SONICS is that it includes the text lyrics of songs, which can aid future research.
@@
Despite the availability of long songs in our dataset, utilizing long-context relationships presents additional challenges. For instance, CNN-based models struggle to capture long-range dependencies due to their inherently local receptive fields. While Transformer-based models can capture these dependencies with global attention, they are computationally expensive with longer audio inputs. To mitigate this trade-off, we introduce Spectro-Temporal Tokens Transformer (SpecTTTra), which uses a Spectro-Temporal Tokenizer to significantly reduce computational costs while employing global attention. Our contributions are summarized as follows:
@@
  • We introduce SONICS, the first large-scale dataset for end-to-end synthetic song detection that addresses the limitations of existing datasets, including limited music-lyrics diversity, short-duration songs, and open fake songs.
  • We provide a human benchmark for fake song detection, filling a gap in previous work, and establish a standard AI benchmark for CNN and Transformer-based models.
@@
  • We propose a faster and memory efficient model, SpecTTTra, which effectively captures long-context temporal relationships in songs while outperforming popular methods.
Table 1: Comparison of Proposed and Existing Fake Song Datasets
@@
CtrSVDD (Zang et al., 2024a) Multi (no English) 4.87 14 164 32,312 188,486 220,798 307
@@ 2 Related Works
Synthetic Speech Detection: The domain of synthetic speech detection, closely tied to synthetic song detection through their shared audio modality, has been extensively explored due to advancements in voice conversion (Zhao et al., 2020) and synthesis techniques (Wang et al., 2021). These advancements have spurred the development of audio spoofing attacks on speaker verification systems and deepfake audio targeting human listeners (Kawa et al., 2023). Synthetic speech detection methods include Light CNN (LCNN) with Max-Feature-Map activations (Lavrentyeva et al., 2019), Transformer encoders with ResNet architectures (Zhang et al., 2021b), RawNet2 with sinc layers and GRU blocks (Tak et al., 2021), and heterogeneous graph attention networks (Jung et al., 2022). However, the unique complexities of songs—such as rhythm, melody, and emotional nuance—present challenges that traditional speech detection methods are not equipped to handle, as shown by Xie et al. (2024) and Zang et al. (2024b). Thus, following CtrSVDD (Zang et al., 2024a), we opted not to conduct similar experiments in our study.
@@
Synthetic Song Detection: Synthetic song detection, a relatively newer and more complex challenge, has gained attention recently. In early 2024, SingFake (Zang et al., 2024b) introduced a dataset of counterfeit songs using Singing Voice Conversion (SVC), along with the task of Singing Voice Deepfake Detection (SVDD) and associated model benchmarks. Subsequent work (Xie et al., 2024; Zang et al., 2024a) combined Singing Voice Synthesis (SVS) with SVC to create phoneme-based songs, leading to specialized detection datasets. Methods in this area include convolutional networks for feature extraction followed by classification using graph neural network  (Jung et al., 2022), wav2vec2-based extraction coupled with graph neural networks (Tak et al., 2022), Linear-Frequency Cepstral Coefficients (LFCC) used with ResNet18 models (Zhang et al., 2021a) and combination of music-specific models (MERT) & linguistic models (wav2vec2.0) with targeted augmentations (Chen et al., 2024). However, SVC and SVS-based datasets retain original background music, leading to a detectable “Karaoke effect” artifact. Recent end-to-end fake songs by Suno and Udio, can produce divergent fake songs where all musical components (e.g., background music, styles, and lyrics) can be synthetic, presenting a severe detection challenge. As shown in Table 2, models trained on SingFake dataset, perform poorly when tested on our end-to-end fake songs dataset, with a significant drop in detection performance (F1 score) ranging from 20% to 64%.
Table 2: Performance of SingFake-trained models on SingFake vs SONICS dataset
Model SingFake SONICS
EER $\downarrow$ F1 $\uparrow$ Sens. $\uparrow$ Spec. $\uparrow$ EER $\downarrow$ F1 $\uparrow$ Sens. $\uparrow$ Spec. $\uparrow$
ConvNeXt 0.20 0.86 0.90 0.65 0.38 0.33 0.22 0.88
ViT 0.29 0.84 0.97 0.35 0.49 0.64 0.87 0.16
EfficientViT 0.19 0.88 0.93 0.64 0.50 0.35 0.28 0.71
@@
Long Audio Classification: Songs exhibit long-range temporal patterns, such as repeated verses, rhythms, etc. setting them apart from speech (Albouy et al., 2024). Despite their potential to enhance detection performance, these patterns have been largely overlooked in existing methods (Xie et al., 2024; Zang et al., 2024b; a). Meanwhile, long audio classification remains a relatively less explored area in audio research. Although automatic speech recognition handles long audio data (Koluguri et al., 2024), it struggles with end-to-end processing of extended audio due to its high computational cost, thus often uses sliding window techniques (Gulati et al., 2020; Radford et al., 2023) to manage costs. This further complicates leveraging long-context features for fake song detection.
@@ 3 Methodology
3.1 SONICS Dataset
The development of a modern synthetic song detection system necessitates a dataset that meets several stringent criteria, which are conspicuously absent in existing music datasets. These criteria include: 1) songs where all components—not just vocals—can be AI-generated; 2) song lengths sufficient to capture long-term contextual relationships; 3) a diverse spectrum of music-lyrics combinations; and 4) a quantity of data substantial enough to serve as a generative model benchmark. Addressing these needs, we introduce the SONICS dataset, a comprehensive collection of end-to-end AI-generated songs produced using the latest audio generative models, spanning lengths from 32 to 240 sec and encompassing an extensive array of music-lyrics styles. A detailed comparison of SONICS with existing datasets is presented in Table 1. It clearly illustrates that datasets such as FSD (Xie et al., 2024), SingFake (Zang et al., 2024b), and CtrSVDD (Zang et al., 2024a) fall short of fulfilling all the outlined criteria when juxtaposed with SONICS. Additionally, a comprehensive distribution summary of the SONICS dataset is provided in Table 3.
Table 3: Summary of our proposed SONICS dataset
[…]
Suno v3.5
[…]
Udio 130
[…]
Suno v3
[…]
Suno v2
[…]
Udio 32
[…]
Suno v3.5
[…]
Suno v3
[…]
Suno v2
[…]
Suno v3.5
[…]
Udio 130
[…]
Suno v3
[…]
Suno v2
[…]
Udio 32
[…]
YouTube
[…]
# Train 774 656 0 0 0 4558 0 0 11863 16172 0 0 0 32686 66709
# Test 77 53 206 100 203 396 634 309 1117 1566 2797 1350 3970 13237 26015
# Valid 9 12 37 16 30 67 117 51 196 286 494 258 700 2167 4440
3.1.1 Real Songs
The dataset’s Real songs segment comprises original compositions by human artists. This portion of the dataset was initially compiled by random sampling from the Genius Lyrics Dataset (J., 2023), which provides metadata including lyrics, titles, and artist information. Subsequently, leveraging this metadata, a search was conducted dynamically on YouTube to retrieve the corresponding audio files. This process yielded 48,090 songs performed by 9,096 artists.
3.1.2 Fake Songs
For generating end-to-end fake songs, we utilized the only two currently available audio generative model families: Suno and Udio. Specifically, we used three variants from Suno—Suno v2, v3, and v3.5. Suno v2, the earliest iteration, generates songs up to 80 seconds (sec.) in length. Suno v3 extends this capability to 120 sec. The latest iteration, Suno v3.5, produces songs up to 240 sec. From the Udio family, we used Udio 32 and Udio 130, capable of generating songs of 32 and 130 sec, respectively. For generating and downloading songs from Suno, we utilized the community API (gcui art, 2024). In contrast, for Udio, we manually generated the songs from the website and then downloaded them using community API (Riera, 2024). Subsequently, we used PyAnnote (Bredin & Laurent, 2021) to filter out audio files lacking vocal components, which were predominantly found among the Udio-generated song.
@@ 3.1.2 Fake Songs
The Fake songs are generated with Suno and Udio models by two primary text inputs: 1) lyrics and 2) song-style (e.g., instruments, genre, vocal type) where for each combination of (lyrics, style) input, two songs are generated. Based on the nature of these inputs, the Fake songs are categorized into three distinct groups: i) Full Fake (FF)—featuring both AI-generated lyrics and song-style; ii) Mostly Fake (MF)—where lyrics are AI-generated based on real lyrics features and song-style is derived from real songs; and iii) Half Fake (HF)—where the lyrics are directly sourced from real songs, with song-style also extracted from real songs. To generate FF songs, we curated a rich set of metadata, including 57 broad topics (e.g., friendship, betrayal), 292 specific topics (e.g., star trek, pokemon), 49 music genres (e.g., rock, metal), and 72 moods (e.g., calm, angry). The full list can be found in the Appendix. Random combinations of these elements were used to generate final lyrics and styles via a Large Language Model (LLM). The complete pipeline for FF song generation is illustrated in Fig. 1. Here the prompt generator generates a prompt for later stages by filling in the variable values into the placeholders of a predefined prompt template (details in Appendix H). In contrast, the creation of MF and HF songs involved extracting song-styles from real songs using the multimodal LLM Gemini 1.5 Flash, which uniquely supports audio input. Additionally, for MF songs, lyrics were generated using the GPT-4o LLM, which extracts features from existing real song lyrics to produce new lyrics with a similar distribution, avoiding direct replication of existing content. The pipeline for generating MF songs is detailed in Fig. 2. The HF songs, differing only in the use of real lyrics directly sourced from existing songs, follow a similar process as MF songs. It is noteworthy that, Udio models cannot generate songs for lyrics from real songs; thus, HF songs contain only Suno models. In total, we generated 6,132 HF songs, 40,769 MF songs, and 2,173 FF songs using Suno and Udio models. The prompt templates used in the LLM to extract lyrics features and song styles, as well as to generate lyrics and styles, can be found in the Appendix.
3.2 Model
The architecture of an audio classification model depends on how the input audio is processed for feature extraction. In our benchmark, we use Mel-spectrogram due to its versatile usage and effective performance across various audio processing tasks (Gong et al., 2021; Radford et al., 2023; Rouard et al., 2023; Niizumi et al., 2024; Shul & Choi, 2024). Spectrograms, resembling 2D shape of an image, allow the usage of image classification models for audio classification. Since image classification models are generally categorized into CNN-based and Transformer-based architectures, we use popular models from both categories in our study, such as ConvNeXt (Liu et al., 2022b) and ViT (Dosovitskiy et al., 2020), for benchmarking. Additionally, we employ EfficientViT (Cai et al., 2023), a hybrid model that integrates both Transformer and CNN components. However, these models encounter challenges when dealing with the inherent long-range temporal dependencies in songs. Specifically, ConvNeXt struggles to capture long-range dependencies due to its local receptive field, a characteristic inherited from CNNs. On the other hand, while ViT is capable of capturing long-range dependencies, it incurs significant computational costs as the number of patches/tokens rapidly increases with longer audio inputs. Similarly, EfficientViT, despite its linear global attention, becomes computationally expensive for long audio due to the large number of tokens combined with its multi-scale operations. To address this issue, we introduce SpecTTTra, which utilizes global attention similar to ViT but employs a Spectro-Temporal Tokenizer to reduce the number of tokens considerably, thereby lowering computational costs and enhancing efficiency significantly.
@@ 3.2 Model
3.2.1 Spectro-Temporal Tokens Transformer (SpecTTTra)